Directions: Solve each system of equations by the substitution method.
Problem: 3x-4y=10 & y=x-3
I have no idea how to do this problem :(

Answer by jim_thompson5910(35256)

Start with 3x-4y = 10 and plug in y = x-3. Then solve for x

3x-4y = 10
3x-4y = 10
3x-4(y) = 10
3x-4(x-3) = 10
3x-4(x-3) = 10
3x-4x+12 = 10
-x-12 = 10
-x = 10-12
-x = -2
x = 2

Then use this to find y

y = x-3
y = 2-3
y = -1

So the solutions are x = 2 and y = -1

The solution as an ordered pair is (2, -1)